OpenStudy (anonymous):

What is the general form of the equation of a circle with its center at (-2, 1) and passing through (-4, 1)? x2 + y2 − 4x + 2y + 1 = 0 x2 + y2 + 4x − 2y + 1 = 0 x2 + y2 + 4x − 2y + 9 = 0 x2 − y2 + 2x + y + 1 = 0

OpenStudy (anonymous):

using the equation, (x-h)^2 +(y-k)^2 = r^2 where the center =(h, k) and the point =(x,y) inserting, (x-(-2))^2 + (y- 1)^2 = r^2 (x+2)^2 + (y- 1)^2 = r^2.........................equation (-4 +2)^2 + (1 -1)^2 = r^2 (-2)^2 + (0)^2 =r^2 4 =r^2 (x+2)^2 + (y- 1)^2 = r^2 (x+2)^2 + (y- 1)^2 = 4 expanding x2 + 4x +4 +y2 -2y +1 =4 x2 +y2+4x-2y +1 =0

OpenStudy (anonymous):

Thank you